OpenStudy (bieber896):

can someone help me on substituting and eliminating systems of equations?! I'm so confused

jimthompson5910 (jim_thompson5910):

Post the full problem please

OpenStudy (bieber896):

x=3y-8 5y=19+4x

OpenStudy (bieber896):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

since x is fully isolated, you can replace the x in the second equation with 3y-8 so `5y=19+4x` turns into `5y=19+4(3y-8)`

jimthompson5910 (jim_thompson5910):

do you see how to solve for y?

OpenStudy (bieber896):

5y=19+12y-32??

OpenStudy (bieber896):

but there's two ys

jimthompson5910 (jim_thompson5910):

get all the y terms to one side so subtract 12y from both sides to get the y terms to the left side

OpenStudy (bieber896):

-7y=19+(-32)?

jimthompson5910 (jim_thompson5910):

good, then combine like terms on the right side

OpenStudy (bieber896):

-7y=-13

jimthompson5910 (jim_thompson5910):

what comes next

OpenStudy (bieber896):

divide both sides by -7?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

\[\Large -7y = -13\] \[\Large \frac{-7y}{-7} = \frac{-13}{-7}\] \[\Large y = \frac{13}{7}\] Notice how the two negatives divide to make a positive result

OpenStudy (bieber896):

so y=1.85

jimthompson5910 (jim_thompson5910):

I'd leave it as a fraction because the decimal result is only an approximation

OpenStudy (bieber896):

oh ok

OpenStudy (bieber896):

so then do I plug that in to the next equation?

jimthompson5910 (jim_thompson5910):

yes, you'll plug that into `x=3y-8` to find x

OpenStudy (bieber896):

so x=-2 3/7

jimthompson5910 (jim_thompson5910):

or x = -17/7 so the solution as an ordered pair is \(\LARGE \left( -\frac{17}{7}, \frac{13}{7}\right)\)

OpenStudy (bieber896):

ok thanks so much ! can we just do one more?!

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (bieber896):

y=3x+3 y=2x+1

jimthompson5910 (jim_thompson5910):

y is equal to 3x+3 AND it's also equal to 2x+1

jimthompson5910 (jim_thompson5910):

since y is equal to those two things, we can set the two equal to each other basically we're performing a substitution

jimthompson5910 (jim_thompson5910):

y=2x+1 3x+3 = 2x+1 ... replace y with 2x+1 solve for x to get x = ??

OpenStudy (bieber896):

so do I have to get rid of one of the variables?

OpenStudy (anonymous):

isolate your x

jimthompson5910 (jim_thompson5910):

move the 2x over you do this by subtracting 2x from both sides

OpenStudy (anonymous):

x=___

OpenStudy (bieber896):

1x+3=+1

jimthompson5910 (jim_thompson5910):

now subtract 3 from both sides to move that 3 to the right side

OpenStudy (bieber896):

1x=-2

jimthompson5910 (jim_thompson5910):

which is the same as x = -2

OpenStudy (bieber896):

yup

OpenStudy (bieber896):

then plug it in

jimthompson5910 (jim_thompson5910):

now plug that into either equation to find y

OpenStudy (bieber896):

so the solution is (-2,3)

OpenStudy (bieber896):

or -3 lol my bad

jimthompson5910 (jim_thompson5910):

yes it's (-2,-3)

OpenStudy (bieber896):

thank you both so much for your help @jim_thompson5910 @tekman1298